satellite73:

\[\frac{dy}{dx}-\frac{2y}{x}=\frac{3y^4}{x^2}\]

satellite73:

this is where we are, right?

Vrefela:

๐Ÿ‘

satellite73:

ok now if the y part was not on the right, this would be one like from 2.3, linear

Vrefela:

So whats the difference between bernoulli and linear?

satellite73:

\[\huge \frac{dy}{dx}-\frac{2y}{x}=\frac{3\color{red}{y^4}}{x^2}\]

satellite73:

you gotta get rid of that \(y^4\) to make it linear. It is bernoulli because that \(y^4\) is there

Vrefela:

Does y have to be on both sides as well to make it a bernoulli? Or can it be x = y^4?

satellite73:

\[y'+p(x)y=f(x)\] is linear \[y'+p(x)y=f(x)y^n\]is bernoulli

Vrefela:

Oh! So linear doesnt have any y on the right side?

satellite73:

no it does not

satellite73:

2.3 had definition

Vrefela:

That makes so much more sense.

Vrefela:

This textbook sucks. I actually miss stewarts now lol

satellite73:

so gimmick is to make bernoulli linear by making a substitution

satellite73:

you make the magic substitution \ and by some miracle it goes

Vrefela:

u = y/x ?

satellite73:

\[\huge u=y^{1-n}\]

Vrefela:

Oops

satellite73:

that was for the other type

Vrefela:

Ok

satellite73:

homo genius

satellite73:

so i guess you missed why the magic trick works right?

Vrefela:

Yes

satellite73:

ok we can see what happens by doing it in this example, although 15 was easier (kinda)

Vrefela:

Okie

satellite73:

\[\huge \frac{dy}{dx}-\frac{2y}{x}=\frac{3\color{red}{y^4}}{x^2}\]evidently \(n=4\) so substitution is ...

satellite73:

\(n=4\)so substitution is ?

Vrefela:

Y^-3

satellite73:

yeah \(u=y^{-3}\) first step is to solve that for \(y\) and then calculate \(\frac{dy}{du}\)

Vrefela:

How?

satellite73:

are you asking me how to solve \(u=y^{-3}\) for \(y\)??

Vrefela:

No I thought to solve for yโ€™ itโ€™s the chain rule

satellite73:

no, first solve for y, then compute \(\frac{dy}{du}\) will make your life much easier before we do any subsitution

Vrefela:

Ok

satellite73:

well...

Vrefela:

Idk

satellite73:

jeez i have to do everything \[u=y^{-3}\\ \huge u^{-\frac{1}{3}}=y\]

Vrefela:

I thought so. I just forgot my minus sign

satellite73:

now what is \(\frac{dy}{du}\) ?

Vrefela:

U^(-1/3)

satellite73:

no, that is just y

satellite73:

take the derivative yo

Vrefela:

0?

satellite73:

better yet, call me, this is taking more time than it is worth

satellite73:

\

satellite73:

\

BRAINIAC:

\(y=u^{-\frac{1}{3}}\) \(\frac{dy}{du}=-\frac{1}{3}u^{-\frac{1}{3}-1}\)